A system of equations is given below, 𝑑π‘₯ + 2𝑦 + 3𝑧 = π‘Ž 2π‘₯ + 3𝑦 βˆ’ 𝑑𝑧 = 𝑏 3π‘₯ + 5𝑦 + (𝑑 + 1)𝑧 = 𝑐 Where 𝑑 is an integer and π‘Ž, 𝑏, 𝑐 are real constants. The system does not have a unique solution, but it is consistent. Show that π‘Ž + 𝑏 = 𝑐.


1
Expert's answer
2021-09-29T08:39:03-0400

By Cramer's rule, the system has many solutions if

"\Delta=\begin{vmatrix}
 t & 2&3 \\
 2 & 3&-t\\
3&5&t+1
\end{vmatrix}=t(8t+3)-2\cdot (5t+2)+3=8t^2-7t-1=0"


"t=\frac{7\pm \sqrt{49+32}}{16}"

"t_1=1,t_2=-1\/8"


"a+b=tx+2y+3z+2x+3y-tz=x(t+2)+5y+z(3-t)"

So, when "t=1" , then:

"a+b=3x+5y+2z=c"